The distance (in miles) for which a certain type of the tyre is to last is a
random variable with the following pdf
1
Expert's answer
2020-05-05T20:12:22-0400

"f(x)=\frac{1}{\theta}e^{\frac{-1}{\theta}x}, x>0"